### v1

@@ 1.1 Our Contributions
Computation User: $O(d)$, Server: $O(d)$ User: $O(n^{2}+dn)$, Server: $O(dn^{2})$ User: $O(dn)$, Server: $O(d)$
Table 1: Comparison of CorDP-DME with existing approaches for DP-DME.
1.2 Related Work
LDP-DME:
As the user has the complete control over the perturbations made to their private vectors, LDP-DME assumes the strongest threat model in DP-DME. Optimality results on LDP-DME have been provided in [21] for $\epsilon\leq 1$ (asymptotic) and in [20, 22] for any $\epsilon$, based on the lower bounds provided in [18]. A number of order optimal LDP-DME algorithms and fundamental results on communication constraints have been provided in [23, 24, 25, 26, 27, 25, 28, 29, 30]. CorDP-DME fundamentally differs from LDP by allowing for the privacy mechanisms among different users to be arbitrarily correlated. In fact, the system model introduced in this work is a generalization of additive DME mechanisms with LDP.
Correlated noise in DP:
Our work is closely related to [31], where the server has access to the perturbed vectors of all users in the DME setting, which are correlated. The DP-DME mechanism introduced in [31] uses a noise generation protocol that is based on SecAgg [12], and achieves the same level of utility as CDP-DME for certain values of $\epsilon$ and $\delta$. Our work deviates from this setting in two main aspects: (i) Dropouts and collusion threshold: CorDP-DME designs the correlated noise distributions for any given thresholds on dropouts and colluding users without bounds. In [31], dropouts are not considered and the collusion threshold is bounded ($<n/3$); (ii) Difference in the problem formulation: The noise generation protocol in [31] uses a fixed correlation coefficient of $-\frac{1}{n}$ among the privacy mechanisms used by different users, that specifies achievable pairs of $(\epsilon,\delta)$ for a target MSE. In this work, we determine the optimum noise distribution (including the correlation coefficient) for any given $\epsilon,\delta,n,c,u$ to minimize the MSE for the required level of privacy.
The advantage of using correlated privacy mechanisms in decentralized learning settings with graph-based approaches is studied in [32, 33]. In these references, the users are represented as nodes in a graph with each edge representing a communication link. The neighboring users (that share an edge) add perfectly (anti) correlated noise to their updates in such a way that they cancel out when averaged. An independent noise component is also added to ensure the privacy of the aggregated updates. The problem setting in [32] differs from ours primarily by not considering the effects of dropouts after noise initialization. In reference [33], dropouts are handled in three additional rounds of user-user communications, unlike CorDP-DME which requires no additional rounds. The line of works related to differentially private follow-the-regularized leader (DP-FTRL) [34, 35, 36, 37] investigate the role of correlated noise across iterations in private learning algorithms.
2 Problem Formulation
We consider a distributed mean estimation (DME) setting with $n$ users, each holding an independent $d$-dimensional vector, and a central server that estimates the mean of these vectors while ensuring $(\epsilon,\delta)$-DP. We assume $n$ point-to-point links, one link from each user to the server in the underlying communication network. Let $[1:n]\coloneq\{1,\dotsc,n\}$ denote the set of all users. Each user $i\in[1:n]$ generates a vector $\mathbf{x}_{i}\in\mathbb{B}^{d}$, where $\mathbb{B}^{d}\subset\mathbb{R}^{d}$ is the unit ball, and sends a distorted version
$M(\mathbf{x}_{i})\coloneq\mathbf{x}_{i}+\mathbf{Z}_{i}.$ (1)
to the server, where $\mathbf{Z}_{i}\in\mathbb{R}^{d}$ are (not necessarily independent) random noise variables. We assume a potentially malicious server and allow for up to a given number of colluding users and dropouts. Let $\mathcal{C}\subset[1:n]$, $|\mathcal{C}|\leq c$ denote the set of users colluding with the server, i.e., the server has access to both $\mathbf{x}_{i}$ and $M(\mathbf{x}_{i})$ for $i\in\mathcal{C}$. Moreover, let $\mathcal{F}\subset[1:n]$, $|\mathcal{F}|\geq t$ be the set of responding users at a given time. We assume $c<t$. The set of non-colluding and responding users is denoted by $\mathcal{U}=\mathcal{F}\setminus\mathcal{C}$.
The server decodes the mean of the vectors of the responding users in $\mathcal{U}$ using a linear function $d_{\mathcal{U}}:\mathbb{R}^{d|\mathcal{U}|}\rightarrow\mathbb{R}^{d}$, defined as:
where $\mathcal{U}=\{j_{1},\dotsc,j_{|\mathcal{U}|}\}$, $j_{1}<j_{2}<\dotsc<j_{|\mathcal{U}|}$, and $\alpha_{i}\in\mathbb{R}$, $i\in\mathcal{U}$ are constants (for a given $\mathcal{F}$ and $\mathcal{C}$) that define the decoding function. Ideally, $\hat{\mathbf{S}}_{\mathcal{U}}$ is approximately the mean of the vectors held by responding users.
We consider a secure DME protocol that consists of an offline (data independent) phase and an online (data dependent) phase. In the offline phase, the users exchange information amongst themselves to ensure that user $i$ has random variable $\mathbf{Z}_{i}$ where $(\mathbf{Z}_{1},\mathbf{Z}_{2},\ldots,\mathbf{Z}_{n})$ have a joint distribution $\pi$. In the online phase, user $i$ sends $M(\mathbf{x}_{i})$ to the server as per (1), and the server outputs $\hat{\mathbf{S}}_{\mathcal{U}}$. Let $\mathcal{I}_{j}$ denote all the information used to generate $\mathbf{Z}_{j}$ for $j\in\mathcal{C}$ in the offline phase. In the DME protocol, after both the offline and online phases, the server has access to $\{\mathcal{I}_{j},\mathbf{x}_{j},\mathbf{Z}_{j}\}_{j\in\mathcal{C}}$, and $\{M(\mathbf{x}_{j})\}_{j=1}^{n}$. Next, we define the privacy constraint considered in this work.
Definition 1 ($(\epsilon,\delta)$-DP in DME)
Let $\mathcal{G}_{i}=\{\{M(\mathbf{x}_{j})\}_{j\in[1:n]\setminus\{i\}},\{\mathbf{x}_{j},\mathbf{Z}_{j},\mathcal{I}_{j}\}_{j\in\mathcal{C}}\}$ denote the information observed by the server from all users except user $i$, for any non-colluding user $i\in[1:n]\setminus\mathcal{C}$. For a given $\epsilon\geq 0$ and $\delta\in(0,1)$, a DP-DME scheme ensures $(\epsilon,\delta)$-DP if the following is satisfied.
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|\mathcal{G}_{i})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\mathcal{G}_{i})+\delta,\quad\forall i\in[1:n]\setminus\mathcal{C}$ (3)
for any fixed $\mathbf{x}_{j}\in\mathbb{B}^{d}$, $j\neq i$, $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}\in\mathbb{B}^{d}$ and $\forall\mathcal{A}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-field.
As the privacy constraint in (3) must be satisfied for any fixed $\mathbf{x}_{j}$,$j\neq i$ for each $i$, Definition 1 implies that $(\epsilon,\delta)$-DP of $\mathbf{x}_{i}$ is satisfied even if the server has access to all $\mathbf{x}_{j}$, $j\neq i$, in addition to $M(\mathbf{x}_{j})$,$\forall j$, and all the information from the colluding users. The privacy constraint in Definition 1 encapsulates the fundamental principles of the original differential privacy (DP) framework as follows. Let $\mathcal{D}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}$ and $\mathcal{D}^{\prime}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}^{\prime}$ be two neighboring datasets that only differ in the vector of user $i$ for any $i\in[1:n]\setminus\mathcal{C}$. For a given dataset $\mathcal{D}$, the information that the server observes is denoted by $\mathcal{V}_{\mathcal{D}}=\{\{M(\mathbf{x}_{j})\}_{j=1}^{n},\{\mathbf{x}_{j},\mathbf{Z}_{j},\mathcal{I}_{j}\}_{j\in\mathcal{C}}\}\in\mathcal{J}$, where $\mathcal{J}$ is the domain of $\mathcal{V}_{\mathcal{D}}$. The original DP framework requires that, for any given $(\epsilon,\delta)$,
$\sup_{\mathcal{D}\sim\mathcal{D^{\prime}}}\mathbb{P}(\mathcal{V}_{\mathcal{D}}\in\mathcal{Y})\leq e^{\epsilon}\mathbb{P}(\mathcal{V}_{\mathcal{D}^{\prime}}\in\mathcal{Y})+\delta$ (4)
must be satisfied for all $\mathcal{Y}\subset\mathcal{J}$. By interpreting (4) as $\mathbb{P}\left(\ln|\frac{\mathbb{P}(\mathcal{V}_{\mathcal{D}}\in\mathcal{Y})}{\mathbb{P}(\mathcal{V}_{\mathcal{D^{\prime}}}\in\mathcal{Y})}|\right)\leq\epsilon)\geq 1-\delta$, for all $\mathcal{D}$,$\mathcal{D}^{\prime}$ [38], we see that (4) is equivalent to (3), when satisfied $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$, for all fixed $\{\mathbf{x}_{j}\}_{j\neq i}$ and for all $i\in[1:n]\setminus\mathcal{C}$. The privacy constraint in Definition 1 is a generalization of the privacy constraints used in LDP and distributed DP with SecAgg, which can be used to investigate the privacy guarantees of an adversarial DME setting. The proofs LDP and distributed DP with SecAgg being special cases of the privacy constraint in Definition 1 are given in appendix A.
Definition 2
An $(n,t,c)$-private DME scheme characterized by the encoding and decoding functions in (1)-(2) is defined by the joint distribution of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ that satisfy Definition 1 and the linear decoding functions $d_{\mathcal{U}}:\mathbb{R}^{d|\mathcal{U}|}\rightarrow\mathbb{R}^{d}$, for each subset of responding but non-colluding users $\mathcal{U}\subseteq[1:n]$.
For a given $(\epsilon,\delta)$-DP constraint, the accuracy of an $(n,t,c)$-private DME scheme is measured by the MSE between the mean estimate at the server $\hat{\mathbf{S}}_{\mathcal{U}}$ and the true mean $\mathbf{S}_{\mathcal{U}}=\frac{1}{|\mathcal{U}|}\sum_{i\in{\mathcal{U}}}\mathbf{x}_{i}$.
Definition 3 (MSE of $(n,t,c)$-private DME)
Consider a DME scheme satisfying $(\epsilon,\delta)$-DP with a given joint distribution $\mathcal{D}_{Z}$ of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ and a decoder $d_{\mathcal{U}}$. For any $\mathcal{F}$ and $\mathcal{C}$, the MSE is defined as,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\mathcal{D}_{Z},\bm{\alpha})\triangleq\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}[\|\hat{\mathbf{S}}_{\mathcal{U}}-\mathbf{S}_{\mathcal{U}}\|_{2}^{2}]$ (5)
where $\bm{\alpha}=[\alpha_{j_{1}},\dotsc,\alpha_{j_{|\mathcal{U}|}}]^{T}$. The expectation is over $\mathcal{D}_{Z}$.
The goal of this work is to characterize the minimum MSE of a given $(n,t,c)$-private DME setting with a given $(\epsilon,\delta)$-DP constraint. We navigate the trade-off between MSE and DP constraints (3) by optimizing the distribution of the privacy mechanism $\mathcal{D}_{Z}$ and the linear decoder at the server:***The order of optimizations in (6) is explained as follows. The privacy mechanism $\mathcal{D}_{Z}$ is determined prior to the mean estimation step in (2), and is fixed irrespective of the number of responding users at a given time. Therefore, the problem is formulated to optimize $\mathcal{D}_{Z}$ for the worst case dropouts and colluding users. The decoder on the other hand is used at the time of estimation as shown in (2), and is optimized based on the number of responding users for any fixed $\mathcal{D}_{Z}$. This allows the decoder to make use of the information on the set of responding users at any given time.
@@ 2 Problem Formulation
$\mathrm{MMSE}\triangleq\inf_{\mathcal{D}_{Z}}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\inf_{\bm{\alpha}}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}[\|\hat{\mathbf{S}}_{\mathcal{U}}-\mathbf{S}_{\mathcal{U}}\|_{2}^{2}]$ (6)
3 Correlated Gaussian Mechanism for DP-DME: A Geometric Interpretation

### v2

@@ I-A Our Contributions
TABLE I: Comparison of CorDP-DME with existing approaches for DP-DME.
I-B Related Work
I-B1 DME with LDP
In DME with LDP, each user perturbs their vectors independently to satisfy $(\epsilon,\delta)$-DP. As the user has the complete control over the perturbations made to their private vectors, DME with LDP assumes the strongest threat model in DP-DME. The strong privacy and security guarantees in LDP comes at a large utility cost. The MSE of DME with LDP is $O(n)$ times higher than the MSE of CDP. Optimality results on DME with LDP have been provided in [29, 12, 11], based on the lower bounds derived in [27]. A number of order optimal LDP-DME algorithms and fundamental results on communication constraints have been provided in [30, 31, 32, 26, 33, 34, 35, 36]. CorDP-DME fundamentally differs from LDP by allowing for the privacy mechanisms among different users to be arbitrarily correlated. In fact, the system model introduced in this work is a generalization of additive DME mechanisms with LDP. The lower bounds established in [27] for LDP are not applicable to CorDP-DME, as they do not account for potential correlations between the privacy mechanisms of different users. These bounds are derived under more restrictive assumptions.
Figure 2: System model: Each user sends a perturbed vector $M(\mathbf{x}_{i})$ and the central server decodes the mean through linear decoding, using the uploads of the responding users. We assume that there can be upto $c$ colluding users and $n-t$ dropouts. The server learns all the random variables observed by the colluding users.
I-B2 Correlated noise in DP-DME
The application of correlated noise in DP-DME has been studied in [16, 17, 14, 37, 24, 25]. Our work is closely related to DECOR [25], CAPE [37], and GOPA [24] as they all incorporate the use of correlated Gaussian noise, and achieve MSEs of the same order as CDP in the absence of dropouts. CAPE and DECOR fundamentally differ from ours by not considering the effects of dropouts. GOPA operates similarly to SecAgg-based DP-DME, but it is designed for general graph structures, and functions over the reals while incorporating correlated noise. The dropout handling mechanisms of both GOPA [24] and SecAgg [18] require multiple rounds of communication, incurring substantial overheads. In fact, because it handles general graph structures, GOPA, unlike SecAgg, does not necessarily achieve CDP performance even after multiple rounds of the protocol. CorDP-DME primarily stands out from GOPA [24], SecAgg-based DP-DME [18], and all comparable correlated privacy mechanisms [16, 17, 14, 37, 24, 25], in that it provides a systematic single-round approach to handle dropouts with no additional communications with the remaining users, no additional shared information among users beyond what is required to generate correlated randomness (unlike [24, 18] which require additional information to be shared among users for dropout recovery), no additional computations at the server, and with no significant loss of MSE. To achieve this, CorDP-DME optimizes the joint distribution of the noise variables across all users to minimize the MSE for the worst-case dropouts and collusions, for any given dropout and collusion thresholds. We provide fundamental privacy-utility trade-offs for arbitrary settings with dropouts and colluding users, considering single-round protocols. Additionally, we present converse results that establish the optimum covariance structure, and derive the optimum noise parameters that minimize the MSE (exact optimal) under different dropout/collusion settings, as opposed to aiming for order-optimality.
I-B3 Secure Aggregation (SA)
SA is a multi-party computation (MPC) protocol that enables computation of the sum of data from multiple participants while ensuring that only the aggregate result is revealed, and individual inputs remain private. The first practical protocol, SecAgg [18], relies on pairwise random noise generation between all user pairs, as described in Section I and Appendix H. Subsequent protocols, such as SecAgg+ [19] and LightSecAgg [20], aim to reduce communication and computational overhead.
SecAgg+ modifies SecAgg by connecting each user with only a randomly selected subset of $O(\log n)$ users for pairwise random noise generation. This reduces the communication and computational overheads of SecAgg+ compared to SecAgg at the expense of a $O\left(\frac{1}{n}\right)$ probability of the protocol failing. [^5] This failure probability is separate from the protocol’s termination due to the number of dropouts exceeding a given threshold. [^6] This is explained in Lemmas 3.8-3.9 in [19]. When $\eta,\sigma_{1}\sim O(\log n)$, the probability of termination is $2^{-\eta}=O\left(\frac{1}{n}\right)$. LightSecAgg removes the need for pairwise noise and eliminates the secret-share reconstruction step used in SecAgg and SecAgg+ during dropout recovery, thereby reducing the overheads. However, it requires the transmission of entire $d$-length vectors during initialization and dropout recovery, as opposed to sharing only the random seeds in SecAgg and SecAgg+, which can be challenging when $d$ is large.
A direct comparison between CorDP-DME (zero probability of failure) and DP-DME with SecAgg+ ($O\left(\frac{1}{n}\right)$ probability of failure) highlights that CorDP-DME achieves lower computational overhead at the server and reduced communication costs during the online phase, requiring $O(d)$ communication per user and $O(dn)$ at the server, compared to SecAgg+’s $O(d+\log n)$ per user and $O(dn+n\log n)$ at the server. However, in the offline phase for noise initialization, CorDP-DME incurs higher communication overhead ($O(n^{2})$ at the server and $O(n)$ per user) than SecAgg+ ($O(n\log n)$ at the server and $O(\log n)$ per user). The key advantage of CorDP-DME over all existing secure aggregation protocols is its single-round design, which eliminates user synchronization issues, avoids additional communication and computation from extra rounds, prevents privacy leaks caused by delayed responses, and mitigates the risk of multi-round privacy attacks.
It is also worth noting that a new version of CorDP-DME can be developed by solving a new problem where a certain probability of protocol failure is allowed, to reduce the initialization costs. This version of CorDP-DME can simply (randomly) select only $O(\log n)$ pairwise random seeds per user to achieve lower initialization (offline) costs as in SecAgg+ while also enjoying its lower online costs. This approach requires optimizing the noise correlations (covariance matrix) among users under a sparsity constraint.
Malicious users and multiple servers: Addressing malicious users is a critical challenge in secure aggregation, as they can compromise the aggregate by providing incorrect inputs. Single-server approaches, such as ACORN [38] and [39], utilize cryptographic techniques for input validation, while multi-server methods like PRIO [40] and [41] employ secret sharing in a non-colluding server setup to validate inputs while maintaining user privacy. In this work, we assume users are honest-but-curious and leave input validation as a topic for future research. Multi-server aggregation methods like PRIO [40] and [41] differ from CorDP-DME in following ways. 1) They use multiple servers and require at least one of them to be non-colluding, while CorDP-DME uses a single server with no assumptions on collusion. 2) In [40] and [41], users act independently, with no correlation required among them, ensuring that user dropouts have no impact on the mechanism. In contrast, CorDP-DME utilizes a privacy mechanism based on correlated noise among participating users, making it essential to account for dropouts. 3) The use of secret sharing in [40] and [41] is fundamentally different from that of CorDP-DME: In PRIO [40] and [41], each user’s input is divided into secret shares and distributed among the non-colluding servers. In CorDP-DME, (partially) additive secret shares of random noise terms are distributed among multiple users.
Other studies on multi-server SA, such as [42, 43], investigate optimal communication rates under perfect user-privacy (conditioned on the aggregate). However, these approaches rely on the strong assumption of non-colluding servers, which can be challenging in adversarial scenarios. CorDP-DME avoids this by operating in a single-server.
II Problem Formulation
We consider a distributed mean estimation (DME) setting with $n$ honest-but-curious users, each holding an independent $d$-dimensional vector, and a central server that estimates the mean of these vectors while ensuring $(\epsilon,\delta)$-DP. We assume a centralized communication model where each user is only connected to the server (see Fig. 2). Let $\mathcal{U}_{all}\coloneq\{1,\dotsc,n\}$ denote the set of all users. Each user $i\in\mathcal{U}_{all}$ generates a private vector $\mathbf{x}_{i}\in\mathbb{B}^{d}$, where $\mathbb{B}^{d}\subset\mathbb{R}^{d}$ is the unit ball, [^7] Even though we present the DP-DME analysis for vectors in the unit ball, this is not a necessary condition. The necessary condition is for the vectors to be bounded. The MSE for a case where the vectors lie in a ball of radius $c$ scales as $c^{2}$ times the MSE of the case corresponding to the unit ball. and sends an encoded (distorted) version of $\mathbf{x}_{i}$ given by,
$M(\mathbf{x}_{i})\coloneq\mathbf{x}_{i}+\mathbf{Z}_{i}$ (1)
to the server, where $\mathbf{Z}_{i}\in\mathbb{R}^{d}$ are (not necessarily independent) random noise variables. We assume an honest-but-curious server that attempts to compromise user-privacy, and allow for up to $c$ colluding users and $n-t$ dropouts. Let $\mathcal{U}_{col}\subset\mathcal{U}_{all}$, $|\mathcal{U}_{col}|\leq c$ denote the set of users colluding with the server, i.e., the server has access to all the random variables observed by the users in $\mathcal{U}_{col}$, including $\mathbf{x}_{i}$, $\mathbf{Z}_{i}$ and $M(\mathbf{x}_{i})$ for $i\in\mathcal{U}_{col}$. Let $\mathcal{U}\subseteq\mathcal{U}_{all}$, $|\mathcal{U}|\geq t$ be the set of responding users at a given time. We assume $c<t$. The goal is to obtain the mean of the vectors of the responding users given by $\mathbf{S}_{\mathcal{U}}=\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}$. Based on the encoded vectors received from $\mathcal{U}$, the server estimates the mean using a linear function $d_{\mathcal{U}}:\mathbb{R}^{d|\mathcal{U}|}\rightarrow\mathbb{R}^{d}$, defined as:
where $\mathcal{U}=\{j_{1},\dotsc,j_{|\mathcal{U}|}\}$, $j_{1}<j_{2}<\dotsc<j_{|\mathcal{U}|}$, and $\alpha_{i}\in\mathbb{R}$ for $i\in\mathcal{U}$ are constants that define the decoding function for a given $\mathcal{U}$. We assume that $\alpha_{i}$ for $i\in\mathcal{U}_{all}$ are independent of $\{M(\mathbf{x}_{j})\}_{j\in\mathcal{U}}$. In this work, we analyze single-round DP-DME settings defined by the encoding and decoding steps in (1)-(2).
In the data encoding step in (1), each user $i$ has to first generate the noise terms $\mathbf{Z}_{i}$, and then encode the private vector as $M(\mathbf{x}_{i})$. We refer to the data-independent noise generation step as the offline phase. The data-dependent vector encoding and decoding steps in (1)-(2) belong to the online phase. Depending on the required covariance structure of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$, each user $j\in\mathcal{U}_{all}$ exchanges a set of random variables denoted by $\mathcal{I}_{j}$ with the other users [^8] For example, $\mathcal{I}_{j}$ can be the set of random seeds shared by user $j$ with other users when generating $\mathbf{Z}_{j}$, such that $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ satisfies some required covariance structure. in the offline phase to facilitate the generation of $\mathbf{Z}_{j}$. [^9] This is a common practice in any correlated privacy mechanism such as SecAgg [18, 44] to generate the correlated noise prior to data communication, with no trusted party present. In successive DP-DME scenarios, such as federated learning with iterative user updates, the offline phase may be executed a single time to establish all necessary shared randomness for the entire process. In the online phase, user $i$ sends $M(\mathbf{x}_{i})$ to the server as per (1), and the server outputs $\hat{\mathbf{S}}_{\mathcal{U}}$. In the DP-DME protocol, after both the offline and online phases, the server has access to the encoded vectors of all users $\{M(\mathbf{x}_{j})\}_{j=1}^{n}$ and all the random variables observed by the colluding users $\{\mathcal{I}_{j},\mathbf{x}_{j},\mathbf{Z}_{j}\}_{j\in\mathcal{U}_{col}}$.
Next, we define the privacy constraint considered in this work, which generalizes the privacy definitions used in LDP [29, 12] and distributed DP with SecAgg [14, 16].
Definition 1 (Generalized $(\epsilon,\delta)$-DP for DME)
Let $\mathcal{D}_{i}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}$ and $\mathcal{D}^{\prime}_{i}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}^{\prime}$ be two neighboring datasets that only differ in the vector of user $i$ for any $i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$. Let $\mathcal{G}_{i}=\{\{M(\mathbf{x}_{j})\}_{j\in\mathcal{U}_{all}\setminus\{i\}},\{\mathbf{x}_{j},\mathbf{Z}_{j},\mathcal{I}_{j}\}_{j\in\mathcal{U}_{col}}\}$ denote all the random variables observed by the server from all users except user $i$, for any $i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$. For a given $\epsilon\geq 0$ and $\delta\in(0,1)$, a DP-DME scheme ensures $(\epsilon,\delta)$-DP if the following is satisfied.
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|\mathcal{G}_{i})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\mathcal{G}_{i})+\delta,$ (3)
$\forall\mathcal{D}_{i},\mathcal{D}^{\prime}_{i}$, $\forall i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$ and $\forall\mathcal{A}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-field.
To motivate the privacy definition in (3), we discuss an alternative DP definition that compares the observed distributions from two neighboring datasets (comparing the joint distributions as opposed to the conditional ones, following the original DP framework), and show that the privacy constraint in Definition 1 is stronger. A detailed explanation is included in Appendix A. Definition 1 generalizes the privacy constraints used in DP-DME with LDP [11, 12, 29] and distributed DP with SA [14, 16], and offers a privacy framework to analyze and compare DP-DME mechanisms with arbitrary correlations among users. The proofs of LDP and distributed DP with SA being special cases of Definition 1 are given in Appendix A. In this work, we use the privacy constraint in Definition 1 to perform fair comparisons of arbitrarily correlated DP mechanisms with the two extreme cases of LDP and distributed DP with SA. In addition, the privacy constraint in Definition 1 gives rise to useful geometric interpretations of the privacy-utility trade-offs in DP-DME, as explained in Section III.
Definition 2 ($(n,t,c,\epsilon,\delta)$-DP-DME scheme)
For a system of $n$ users with up to $c$ colluding users and at least $t$ responding users, and given privacy parameters $(\epsilon,\delta)$, a DP-DME scheme is defined by:
  • the encoding mechanism: the joint distribution $\mathcal{D}_{Z}$ of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ that satisfies the gnearlized $(\epsilon,\delta)$-DP for DME as given in Definition 1, and
  • the decoding mechanism: linear decoding functions $d_{\mathcal{U}}:\mathbb{R}^{d|\mathcal{U}|}\rightarrow\mathbb{R}^{d}$, for each subset $\mathcal{U}\subseteq\mathcal{U}_{all}$.
The accuracy of an $(n,t,c,\epsilon,\delta)$-DP-DME scheme is measured by the MSE between the mean estimate at the server $\hat{\mathbf{S}}_{\mathcal{U}}$ and the true mean $\mathbf{S}_{\mathcal{U}}=\frac{1}{|\mathcal{U}|}\sum_{i\in{\mathcal{U}}}\mathbf{x}_{i}$.
Definition 3 (MSE of an $(n,t,c,\epsilon,\delta)$-DP-DME scheme)
Consider an $(n,t,c,\epsilon,\delta)$ DP-DME scheme with a joint distribution $\mathcal{D}_{Z}$ of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ and a decoder $d_{\mathcal{U}}$ characterized by $\boldsymbol{\alpha}_{\mathcal{U}}=[\alpha_{j_{1}},\dotsc,\alpha_{j_{|\mathcal{U}|}}]^{T}$ for each $\mathcal{U}\subseteq\mathcal{U}_{all}$. For a given $\mathcal{U}\subseteq\mathcal{U}_{all}$, the MSE is defined as,
$\mathrm{MSE}(\mathcal{D}_{Z},\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
$\triangleq\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\mathbb{E}\left[\left\|\hat{\mathbf{S}}_{\mathcal{U}}-\mathbf{S}_{\mathcal{U}}\right\|^{2}\right]$ (4)
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}M(\mathbf{x}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in{\mathcal{U}}}\mathbf{x}_{i}\right\|^{2}\right]$ (5)
@@ II Problem Formulation
The goal of this work is to find the optimum $(n,t,c,\epsilon,\delta)$-DP-DME scheme that minimizes the MSE in Definition 3 for the worst case dropouts and colluding users. Specifically, we characterize the minimum MSE given by,
$\mathrm{MMSE}$
@@
$\triangleq\inf_{\mathcal{D}_{Z}}\quad\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\quad\inf_{\boldsymbol{\alpha}_{\mathcal{U}}}\quad\mathrm{MSE}(\mathcal{D}_{Z},\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$ (6)
$=\inf_{\mathcal{D}_{Z}}\quad\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\quad\inf_{\boldsymbol{\alpha}_{\mathcal{U}}}$
$\quad\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}M(\mathbf{x}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in{\mathcal{U}}}\mathbf{x}_{i}\right\|^{2}\right]$ (7)
while satisfying the privacy constraint in Definition 1 for any $\mathcal{U}_{col}\subset\mathcal{U}_{all}$ with $|\mathcal{U}_{col}|\leq c$. The parameters to be optimized in (7) are the joint distribution $\mathcal{D}_{Z}$ of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ considering the worst case dropouts and colluding users, and the decoder $\boldsymbol{\alpha}_{\mathcal{U}}$ for each $\mathcal{U}\subseteq\mathcal{U}_{all}$. [^10] We show that the optimum $\alpha_{\mathcal{U}}$ only depends on $|\mathcal{U}|$ and not directly on each $\mathcal{U}$ (See Proposition 1 for details). The $\max$ and $\sup$ terms in (7) characterize the worst case dropouts and the private vectors of the users, that give the worst case MSE. In this paper, we establish a minimax result by minimizing the maximum MSE. The order of optimizations in (7) is explained as follows. The noise distribution $\mathcal{D}_{Z}$ is determined prior to the mean estimation step in (2), and is fixed irrespective of the number of responding users at a given time. Therefore, the problem is formulated to optimize $\mathcal{D}_{Z}$ for the worst case dropouts and colluding users. The decoder on the other hand is used at the time of estimation as shown in (2), and is optimized based on the number of responding users for any fixed $\mathcal{D}_{Z}$. This allows the decoder to make use of the information on the set of responding users at any given time.
III Correlated Gaussian Mechanism for DP-DME: A Geometric Interpretation